OpenStudy (anonymous):

Let f(x) = x – 6 and g(x) = x2 – x + 4. Find g(f(x)).

OpenStudy (anonymous):

I'm not sure what to plug in where...

OpenStudy (mathmate):

Let f(x) = x – 6 and g(x) = x2 – x + 4. Find g(f(x)). g(f(x)) = g(x – 6) = (x-6)^2 - (x-6) + 4 = x^2-12x+36 -x +6 +4 = x^2 -13x + 46

OpenStudy (anonymous):

Awesome. Thanks!

OpenStudy (mathmate):

You're welcome! :)